OpenStudy (anonymous):

The formula below can be used to convert temperatures in degrees Fahrenheit, F, to temperatures in degrees Celsius, C. C = 5/9 (F − 32) On a certain day temperatures at the North Pole were between −20°F and −15°F. Which of these is a reasonable temperature in degrees Celsius for that day at the North Pole? F −30°C G −22°C H −27°C J −11°C

OpenStudy (anonymous):

it's -27 according to my calculations...

OpenStudy (anonymous):

first i i took the function to be C(f)=5/9(f-32) then made my domain \[-20\le F \le -15\]

OpenStudy (anonymous):

i checked the values, i got: C(-20)=-28.889 and C(-15)=-26.111 The only value that they give which lies between -28.889 and -26.111, is H: -27.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome! :)